# Advanced Payment Security System: XGBoost, LightGBM and SMOTE Integrated

### v2

@@ I Introduction
Across the globe, network payment platforms are emerging as a burgeoning industry. With the widespread adoption of online payments[6], these platforms have brought about profound and extensive changes to society, fundamentally altering our lifestyles. Network payments involve issues related to fund security and are closely linked to societal stability[54, 23]. The problems inherent in online payments have long constrained the development of e-commerce in the modern financial sector. Various phenomena such as transaction fraud and payment fraud frequently occur, causing significant distress for both bankers and users[44, 46], and represent one of the primary obstacles to the development of network payments[26].
To address these issues, predecessors have summarized and explored various powerful and reliable classification algorithms, including Neural Networks, Logistic Regression, and SVM.[39, 41]. However, these algorithms still face challenges such as complex hyperparameter tuning, handling multi-class problems, long training times, and unsuitability for large-scale datasets. The newly introduced LightGBM and XGBoost have made significant improvements in large-scale data processing, nonlinear handling, training speed, and performance tuning. Nonetheless, they still have drawbacks in handling nonlinear relationships and susceptibility to outliers. Combining their strengths for comprehensive optimization remains a highly challenging task.
To build on the foundation of these models, we plan to combine various technologies and models to explore more powerful and effective solutions. In this article, we will focus on some of the most commonly used models in banking today [12, 32]. After processing fraud data features, we will compare XGBoost, LightGBM, CatBoost, and traditional models. We will then attempt to combine these models and further optimize them using SMOTE [34]. Through detailed experimental comparisons of parameters such as F1 score, precision, recall, and AUC, we aim to develop a new model.
@@ II Related work
Sayjadah [35] has implemented machine learning models to build a robust payment security system to prevent fraud in transactions. He pioneered the use of logistic regression, decision trees, and random forests, achieving an accuracy rate of 82% in risk prediction [27]. While his work laid a solid foundation for our research direction, it also has limitations such as poor hyperparameter performance, unsuitability for large-scale data processing, and inadequate performance in handling multi-class classification problems.
Chen and Guestrin (2016) [3] provided a detailed explanation of XGBoost’s principles and training characteristics. Their research has significantly influenced our work, as we adopted many of their methodologies to enhance our results. XGBoost’s efficiency and performance improvements make it an essential component of our model ensemble. LightGBM (Light Gradient Boosting Machine), developed by Ke et al. (2017) [14], is another powerful gradient boosting framework designed for speed and efficiency. We also integrated LightGBM into our study due to its ability to handle large-scale data with high efficiency, making it a crucial model in our experiments.Although both models are excellent, they also have their own drawbacks and issues, such as overfitting, sensitivity to outliers, and training time.
For our ensemble model, we reviewed various models that matched our requirements. Ultimately, inspired by Seewald [36], we decided to build our new classifier using the stacking approach. Seewald’s work on improving the effectiveness and efficiency of stacking provided a strong foundation for our ensemble method.
The article ”A Survey on Applications of Artificial Intelligence for Smart Manufacturing: Real-Time Production Quality Improvement and System Optimization” offers significant insights that can be directly applied to the development of advanced payment protection systems [2]. This survey highlights how artificial intelligence (AI) techniques [26, 56], such as machine learning, deep learning, and reinforcement learning, are revolutionizing manufacturing processes by improving quality and optimizing system performance.The aforementioned article serves as an important reference for us in model construction and processing.
Outlier detection and handling is an indispensable component of data preprocessing [8, 33]. By undertaking this procedure, we aimed to ensure the quality and integrity of the data, enhance the performance of subsequent machine learning models, uncover valuable insights embedded within the data[31], mitigate the business impact of anomalies, and satisfy relevant regulatory and compliance requirements.
@@ III-A Payment Security System Architecture
As shown in Figure 1, our system is based on accumulated historical transaction data. Depending on different applications and previous experimental results[4, 30], techniques such as Logistic Regression and Neural Networks have already been widely applied in similar system models. These models perform an initial classification of new transaction requests[12], approving legitimate payments and forwarding potential fraud transactions to the Fraud Protection Service for further processing. Therefore, the accuracy of the models is crucial to the system’s performance.
In this research, we will use a series of techniques including XGBoost, LightGBM, and CatBoost to process the data and perform comparisons to identify the best-performing model[11]. Subsequently, we will employ the stacking ensemble technique to construct a new, more powerful ensemble model[43, 42]. This model, combined with SMOTE (Synthetic Minority Over-sampling Technique), will be compared against a series of classic models for research and validation.
@@ III-B Data Processing
In our data processing phase, the primary objectives were to clean missing values and outliers[17], as well as to select the most relevant features for training. This preprocessing is crucial to enhance the effectiveness of our model training[5]. The steps involved in our data processing are outlined below:
@@ III-B3 Dimensionality reduction
In this data processing task, we employed the T-SNE technique to perform dimensionality reduction on high-dimensional data, achieving remarkable results[9]. T-SNE uses non-linear mapping to project data points from a high-dimensional space to a low-dimensional space while preserving the similarity structure between data points as much as possible[47, 20]. This allows us to visually observe the intrinsic patterns and clustering structure of the data in a two-dimensional or three-dimensional space, thereby better understanding the characteristics and distribution of the data. Dimensionality reduction brings multiple benefits. it significantly reduces computational complexity, eliminates redundant information, solve the data sparse problem[25].
@@ III-B4 Synthetic Minority Over-sampling Technique (SMOTE)
When dealing with imbalanced datasets, classifiers often tend to favor the majority class, overlooking the minority class samples. This bias significantly reduces the model’s performance on the minority class, which can be critical in scenarios where minority class detection is essential. To address this issue, this study employs the Synthetic Minority Over-sampling Technique (SMOTE) to balance the dataset[1, 36].
@@ IV-B Experiement Result
In this section, we conduct a detailed comparison of various models used in a Fraud Protection System. To ensure the reliability of our model comparison, we not only employed traditional models such as Logistic Regression and Decision Tree but also incorporated some of the most popular high-performance models available today[53]. This includes deep learning-based models like TabNet, CatBoost, and standalone models such as LightGBM and XGBoost[38]. This comprehensive comparison aims to evaluate the detailed performance of each model while also assessing their practical utility in modern-day scenarios.
@@
By comparing the above parameters and images, we can see that the precision of the Ensemble Model is 0.98, the highest among all models[45, 48], tied with the Support Vector Classifier. This means that among all samples predicted as positive, the Ensemble Model has the highest correct prediction rate and the lowest false positive rate. Additionally, the AUC score of the Ensemble Model is 0.99, the highest among all models. The higher the AUC score, the stronger the model’s performance to identify positive and negative samples. The Ensemble Model performs best in this regard, indicating its superior overall classification ability[21].
Even in terms of Recall and F1 Score, the recall rate of the Ensemble Model is 0.90, which, although not the highest, is not significantly different from other models. The slightly lower recall rate could be because the Ensemble Model emphasizes precision more[50], but it still maintains a high overall level. The F1 Score balances Precision and Recall, with the Ensemble Model achieving an F1 Score of 0.94, consistent with Logistic Regression, XGBoost, LightGBM, CatBoost, and Neural Network. This indicates that the Ensemble Model performs excellently in balancing precision and recall[28, 19].
As an ensemble model that combines multiple base models, the Ensemble Model integrates the prediction results of several base models (such as Logistic Regression, XGBoost, Neural Network, etc.)[55, 7], fully leveraging the strengths of each model and compensating for the shortcomings of single models. Because it aggregates the results of multiple models, the Ensemble Model has better resistance to overfitting and underfitting issues of individual models, making the overall model more stable. By combining multiple models, the Ensemble Model can better capture the diversity of data distributions, improving generalization to new data.
In summary, the Ensemble model combined with SMOTE demonstrates outstanding performance across various metrics[52], particularly excelling in precision and AUC value, highlighting its superiority in practical applications[16, 40].
@@ Conclusion
In today’s society, where transaction fraud poses an escalating threat, the development of robust Payment Security Systems is paramount[51, 24], . This research is driven by the critical need to combat this evolving challenge. Preventing fraud demands not only vast amounts of data but also sophisticated models capable of accurate predictions. Our study rigorously compared several cutting-edge models[22, 29], leveraging the power of SMOTE to further enhance their performance.

### v3

@@ I Introduction
Across the globe, network payment platforms are emerging as a burgeoning industry. With the widespread adoption of online payments[3], these platforms have brought about profound and extensive changes to society, fundamentally altering our lifestyles. Network payments involve issues related to fund security and are closely linked to societal stability. The problems inherent in online payments have long constrained the development of e-commerce in the modern financial sector. Various phenomena such as transaction fraud and payment fraud frequently occur, causing significant distress for both bankers and users, and represent one of the primary obstacles to the development of network payments.
To address these issues, predecessors have summarized and explored various powerful and reliable classification algorithms, including Neural Networks, Logistic Regression, and SVM. However, these algorithms still face challenges such as complex hyperparameter tuning, handling multi-class problems, long training times, and unsuitability for large-scale datasets. The newly introduced LightGBM and XGBoost have made significant improvements in large-scale data processing, nonlinear handling, training speed, and performance tuning. Nonetheless, they still have drawbacks in handling nonlinear relationships and susceptibility to outliers. Combining their strengths for comprehensive optimization remains a highly challenging task.
To build on the foundation of these models, we plan to combine various technologies and models to explore more powerful and effective solutions. In this article, we will focus on some of the most commonly used models in banking today . After processing fraud data features, we will compare XGBoost, LightGBM, CatBoost, and traditional models. We will then attempt to combine these models and further optimize them using SMOTE. Through detailed experimental comparisons of parameters such as F1 score, precision, recall, and AUC, we aim to develop a new model.
@@ II Related work
Sayjadah has implemented machine learning models to build a robust payment security system to prevent fraud in transactions. He pioneered the use of logistic regression, decision trees, and random forests, achieving an accuracy rate of 82% in risk prediction. While his work laid a solid foundation for our research direction, it also has limitations such as poor hyperparameter performance, unsuitability for large-scale data processing, and inadequate performance in handling multi-class classification problems.
Chen and Guestrin (2016) provided a detailed explanation of XGBoost’s principles and training characteristics. Their research has significantly influenced our work, as we adopted many of their methodologies to enhance our results. XGBoost’s efficiency and performance improvements make it an essential component of our model ensemble. LightGBM (Light Gradient Boosting Machine), developed by Ke et al. (2017) , is another powerful gradient boosting framework designed for speed and efficiency. We also integrated LightGBM into our study due to its ability to handle large-scale data with high efficiency, making it a crucial model in our experiments.Although both models are excellent, they also have their own drawbacks and issues, such as overfitting, sensitivity to outliers, and training time.
For our ensemble model, we reviewed various models that matched our requirements. Ultimately, inspired by Seewald, we decided to build our new classifier using the stacking approach. Seewald’s work on improving the effectiveness and efficiency of stacking provided a strong foundation for our ensemble method.
The article ”A Survey on Applications of Artificial Intelligence for Smart Manufacturing: Real-Time Production Quality Improvement and System Optimization” offers significant insights that can be directly applied to the development of advanced payment protection systems [1]. This survey highlights how artificial intelligence (AI) techniques , such as machine learning, deep learning, and reinforcement learning, are revolutionizing manufacturing processes by improving quality and optimizing system performance.The aforementioned article serves as an important reference for us in model construction and processing.
Outlier detection and handling is an indispensable component of data preprocessing . By undertaking this procedure, we aimed to ensure the quality and integrity of the data, enhance the performance of subsequent machine learning models, uncover valuable insights embedded within the data[12], mitigate the business impact of anomalies, and satisfy relevant regulatory and compliance requirements.
@@ III-A Payment Security System Architecture
As shown in Figure 1, our system is based on accumulated historical transaction data. Depending on different applications and previous experimental results, techniques such as Logistic Regression and Neural Networks have already been widely applied in similar system models. These models perform an initial classification of new transaction requests[6], approving legitimate payments and forwarding potential fraud transactions to the Fraud Protection Service for further processing. Therefore, the accuracy of the models is crucial to the system’s performance.
In this research, we will use a series of techniques including XGBoost, LightGBM, and CatBoost to process the data and perform comparisons to identify the best-performing model[5]. Subsequently, we will employ the stacking ensemble technique to construct a new, more powerful ensemble model. This model, combined with SMOTE (Synthetic Minority Over-sampling Technique), will be compared against a series of classic models for research and validation.
@@ III-B Data Processing
In our data processing phase, the primary objectives were to clean missing values and outliers, as well as to select the most relevant features for training. This preprocessing is crucial to enhance the effectiveness of our model training. The steps involved in our data processing are outlined below:
@@ III-B3 Dimensionality reduction
In this data processing task, we employed the T-SNE technique to perform dimensionality reduction on high-dimensional data, achieving remarkable results. T-SNE uses non-linear mapping to project data points from a high-dimensional space to a low-dimensional space while preserving the similarity structure between data points as much as possible. This allows us to visually observe the intrinsic patterns and clustering structure of the data in a two-dimensional or three-dimensional space, thereby better understanding the characteristics and distribution of the data. Dimensionality reduction brings multiple benefits. it significantly reduces computational complexity, eliminates redundant information, solve the data sparse problem.
@@ III-B4 Synthetic Minority Over-sampling Technique (SMOTE)
When dealing with imbalanced datasets, classifiers often tend to favor the majority class, overlooking the minority class samples. This bias significantly reduces the model’s performance on the minority class, which can be critical in scenarios where minority class detection is essential. To address this issue, this study employs the Synthetic Minority Over-sampling Technique (SMOTE) to balance the dataset.
@@ IV-B Experiement Result
In this section, we conduct a detailed comparison of various models used in a Fraud Protection System. To ensure the reliability of our model comparison, we not only employed traditional models such as Logistic Regression and Decision Tree but also incorporated some of the most popular high-performance models available today. This includes deep learning-based models like TabNet, CatBoost, and standalone models such as LightGBM and XGBoost. This comprehensive comparison aims to evaluate the detailed performance of each model while also assessing their practical utility in modern-day scenarios.
@@
By comparing the above parameters and images, we can see that the precision of the Ensemble Model is 0.98, the highest among all models, tied with the Support Vector Classifier. This means that among all samples predicted as positive, the Ensemble Model has the highest correct prediction rate and the lowest false positive rate. Additionally, the AUC score of the Ensemble Model is 0.99, the highest among all models. The higher the AUC score, the stronger the model’s performance to identify positive and negative samples. The Ensemble Model performs best in this regard, indicating its superior overall classification ability.
Even in terms of Recall and F1 Score, the recall rate of the Ensemble Model is 0.90, which, although not the highest, is not significantly different from other models. The slightly lower recall rate could be because the Ensemble Model emphasizes precision more, but it still maintains a high overall level. The F1 Score balances Precision and Recall, with the Ensemble Model achieving an F1 Score of 0.94, consistent with Logistic Regression, XGBoost, LightGBM, CatBoost, and Neural Network. This indicates that the Ensemble Model performs excellently in balancing precision and recall.
As an ensemble model that combines multiple base models, the Ensemble Model integrates the prediction results of several base models (such as Logistic Regression, XGBoost, Neural Network, etc.), fully leveraging the strengths of each model and compensating for the shortcomings of single models. Because it aggregates the results of multiple models, the Ensemble Model has better resistance to overfitting and underfitting issues of individual models, making the overall model more stable. By combining multiple models, the Ensemble Model can better capture the diversity of data distributions, improving generalization to new data.
In summary, the Ensemble model combined with SMOTE demonstrates outstanding performance across various metrics, particularly excelling in precision and AUC value, highlighting its superiority in practical applications.
@@ Conclusion
In today’s society, where transaction fraud poses an escalating threat, the development of robust Payment Security Systems is paramount, . This research is driven by the critical need to combat this evolving challenge. Preventing fraud demands not only vast amounts of data but also sophisticated models capable of accurate predictions. Our study rigorously compared several cutting-edge models, leveraging the power of SMOTE to further enhance their performance.